A runner goes around a band on a track with a radius of 52 m at a speed of 2.0 m/s. Determine the centripetal acceleration of the runner.
1
Expert's answer
2020-04-13T10:10:05-0400

"a=\frac{v^2}{r}"

"a=\frac{2^2}{52}=0.077\frac{m}{s^2}"